Three cells each with an emf of 2,5 V and an internal resistance of 0,2 ohms are connected in series and across 15 ohms

1
Expert's answer
2022-02-15T19:05:02-0500

Total emf = 2.5+2.5+2.5=7.5 volt

Total resistance of cells =0.2+0.2+0.2=0.6 ohm

Total resistance of circuit= 15+.6=15.6 ohm

Current in circuit=7.5/15.6=0.48 A